OpenStudy (anonymous):

Frank’s quarter grades are 82.5, 94.7, and 87.9. What grade, rounded to the nearest percent, will Frank have to earn in the last quarter to receive at least a 88% final course grade? A: 91% , B: 87%, C: 95%, D: 96%

OpenStudy (anonymous):

I would try substitution for this problem. The average for 82.5, 94.7, and 87.9 with a 95 C answer is a 90% Try something lower than that.

OpenStudy (anonymous):

91 averaged in with the other two numbers is 88% From looking at the choices we can evaluate that the answer is B 87%

OpenStudy (anonymous):

Does that make sense? I went a little fast.

OpenStudy (texaschic101):

(82.5 + 94.7 + 87.9 + x) / 4 >= 88 (265.1 + x) / 4 > = 88 -- multiply by 4 265.1 + x >= (88)(4) 265.1 + x >= 352 -- subtract 265.1 x >= 352 - 265.1 x >= 86.9 rounds to 87

OpenStudy (mathmale):

I'd suggest going back to the definition of mean or average, and choosing a variable (such as x) to represent the unknown grade. In this case, the average, which is 88, is derived from (sum of all the grades)/(number of grades): 88 = (82.5+94.7+87.9+x)/4. Solve this for x.